OpenStudy (ray1998):

Fan and medal!!! Simplify the expression. (-2i)(8i) Thank you in advance! :)

OpenStudy (anonymous):

do u know what i^2 stands for

OpenStudy (anonymous):

so ur first step should be multiplying i

OpenStudy (anonymous):

\[-2i \times 8i= -2i^2 8\]

OpenStudy (anonymous):

r u getting it?

OpenStudy (anonymous):

because i^2= -1

OpenStudy (anonymous):

so we substitute that an we get \[-2 \times -1 \times 8=\]

OpenStudy (anonymous):

16

OpenStudy (ray1998):

I'M SO SORRY! I've been very busy lately. Thank you so much for replying. And I'm very sorry to have answered so late. @ayeshaafzal221

OpenStudy (anonymous):

thats perfectly fine let me know if u need more help just tag me:)

OpenStudy (ray1998):

I will :) thank you very much for you time and effort! @ayeshaafzal221